OpenStudy (anonymous):

Graph the circle (x+2)2 + y 2 = 16. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

(x+2)^2 + y^2 = 16 is the same as (x-(-2))^2 + (y-0)^2 = 4^2 so are you able to determine the center and radius?

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

2 is what

OpenStudy (anonymous):

the radius

jimthompson5910 (jim_thompson5910):

no, but close

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

compare (x-(-2))^2 + (y-0)^2 = 4^2 to (x-h)^2 + (y-k)^2 = r^2

jimthompson5910 (jim_thompson5910):

you'll see that r^2 = 16 ----> r = 4

jimthompson5910 (jim_thompson5910):

so radius is 4

jimthompson5910 (jim_thompson5910):

yep, good you got it the second time

OpenStudy (anonymous):

and where should my center be?

jimthompson5910 (jim_thompson5910):

the center is what

jimthompson5910 (jim_thompson5910):

well (h,k) is your center, so ???

jimthompson5910 (jim_thompson5910):

if you can find h and k, then you can find the center

OpenStudy (anonymous):

2,2?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

again compare (x-(-2))^2 + (y-0)^2 = 4^2 to (x-h)^2 + (y-k)^2 = r^2

jimthompson5910 (jim_thompson5910):

h = ??? k = ???

OpenStudy (anonymous):

hmm im not sure what to do there

jimthompson5910 (jim_thompson5910):

notice how the -(-2) part matches up with the -h part

jimthompson5910 (jim_thompson5910):

so h = -2

jimthompson5910 (jim_thompson5910):

y-0 matches up with y-k so k = 0

jimthompson5910 (jim_thompson5910):

(h,k) = (-2,0)

jimthompson5910 (jim_thompson5910):

your center is (-2,0)

OpenStudy (anonymous):

ohh

jimthompson5910 (jim_thompson5910):

That's enough info to graph

OpenStudy (anonymous):

ok thanks :)

jimthompson5910 (jim_thompson5910):

yw